OpenStudy (freemap):

Could Some please check my answer? The quotient of a number and 5 is equal to 30. What is the number? 150 is my answer

OpenStudy (anonymous):

yes you are right cause 150 divide by 5 is 30

OpenStudy (anonymous):

or you could do it another way

OpenStudy (freemap):

thank could you help me with another one?

OpenStudy (anonymous):

sure

OpenStudy (freemap):

Which equation represents the sentence? Four more than one-third a number is equal to 14.

OpenStudy (freemap):

getting the options

OpenStudy (anonymous):

ok np

OpenStudy (freemap):

A. 4n+1/3=14 B. 4(n/3)=14 C. n/3+4=14 D. 1/3(n+4)=14

OpenStudy (anonymous):

which one do you think it is?

OpenStudy (freemap):

D

OpenStudy (freemap):

c

OpenStudy (anonymous):

are you sure?

OpenStudy (freemap):

yes c

OpenStudy (anonymous):

you are correct

OpenStudy (freemap):

Thanks you

OpenStudy (anonymous):

np

OpenStudy (freemap):

Please one more

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do you mind helping me when you done?

OpenStudy (freemap):

A photograph has a width of 8 cm and a length of 3x cm. Which equation represents the area A of the photograph? A. =11x A. =24x My answer A. =x+24 A= 3x+8

OpenStudy (freemap):

Yes I'll try to help

OpenStudy (anonymous):

yes you are correct nice job!

OpenStudy (freemap):

Thanks how can i help you

OpenStudy (anonymous):

find my question first in the post

OpenStudy (freemap):

ok